Circular No. 8941

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 J2 (BESHORE) E. Beshore reports his discovery of a comet with a bright, condensed

condensation ∼ 20′′ across and a faint tail in p.a. 270^{o} on images taken with the Mt. Lemmon 1.5-m reflector (discovery observation tabulated below). After posting on the Minor Planet Center’s ‘NEOCP’ webpage, J. C. Pelle reported a diffuse head and a tail in p.a. 250^{o} on images taken by N. Teamo (Punaauia, French Polynesia, 0.41-m reflector) on May 6.52–6.56 UT. A. C. Gilmore (Mt. John 1.0-m reflector, May 6.53–6.55) found a bright circular coma of diameter ∼ 16′′ (total mag 14.4–14.5 in a 7′′-radius aperture) and a faint, short tail seen as slight diffuseness on the west side (p.a. ∼ 250^{o}). G. Sostero and E. Guido (Remanzacco, Italy, 0.2-m reflector, May 6.95) write that co-added unfiltered exposures show an extremely compact 20′′ coma of almost uniform brightness, slightly elongated toward the southwest; remote Cousins-R-band imaging with a 0.25-m reflector near Mayhill, NM, U.S.A., on May 7.37 yields coma diameter ∼ 25′′ (total mag ∼ 13.9), elongated toward p.a. 240^{o}. Other reports of cometary appearance have been received from P. Birtwhistle, Great Shefford, Berkshire, England (0.40-m reflector, May 6.96–6.97; diffuse with a concentrated center, the 22′′ coma having a softer edge towards p.a. 260^{o} but no obvious tail); J. M. Aymami, Tiana, Spain (0.25-m reflector, May 6.97–6.98; nebulous object of mag ∼ 14.8; L. Buzzi, Varese, Italy (0.60-m reflector, May 7.03–7.09; several stacked images reveal a compact coma, clearly elongated with a broad tail at least 20′′ long); C. Jacques and E. Pimentel, Belo Horizonte, Minas Gerais, Brazil (0.30-m reflector, May 7.08–7.09; condensed, elongated coma of size 18′′.7 × 31′′.3); J. E. McGaha, Tucson, AZ, U.S.A. (0.36-m reflector, May 7.29– 7.32; starlike nuclear condensation with a 3′′ coma, elongated 5′′ toward p.a. 260^{o}); and J. Young, Table Mountain (0.61-m reflector, May 7.32–7.35; 12′′ coma, slightly elongated toward p.a. 260^{o}, with a hint of a poorly defined tail ≥ 40′′ long between between p.a. 245^{o} and 275^{o}).

2008 UT  α_{2000}  δ_{2000}  Mag.
May 6.44680  17^{h}06^{m}21.^{s}38  −13^{o}33′36′′.1  13.8

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-J39; this comet may be of short period.

T = 2008 Oct. 7.073 TT  ω = 256.^{o}546
Ω = 85.219 }2000.0
q = 1.09077 AU  i = 16.563
2008 May 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green